OpenStudy (anonymous):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? a) .2 seconds to 1.3 seconds b) .6 seconds to 1.8 seconds c) .5 seconds to 2.1 seconds d) .4 seconds to 1.9 seconds

OpenStudy (rogue):

Let the height/position function equal to 9 and solve for t like a normal quadratic equation. The ball will be above 9 feet between those two t values.

OpenStudy (anonymous):

but during what time interval is the height of the ball above 9 feet? (and you have to Round the answer to the nearest tenth of a second.)

OpenStudy (rogue):

The time interval is the interval between the solutions to h(t) = 9.

OpenStudy (anonymous):

so which answer choice would best fit this problem?

OpenStudy (callisto):

Put \(h(t) \ge 9\) \(-16t^2 + 24t + 5 \ge 9 \) \(-16t^2 + 24t -4 \ge 0 \) \(4t^2 -6 t +1 \ge 0 \) Can you solve it?

OpenStudy (anonymous):

no -.- i suck in math

OpenStudy (callisto):

Do you know how to solve a quadratic equation?

OpenStudy (callisto):

Sorry, the last line there should be \(4t^2 -6 t +1 \le 0\) <- solve this

OpenStudy (anonymous):

is this right?

OpenStudy (callisto):

Use this formula to solve -> yes But the range should be like \(x_1\le x \le x_2\) where \(x_1\) and \(x_2\) are the values you get from that formula.

OpenStudy (anonymous):

so is it a,b,c, or d ? a) .2 seconds to 1.3 seconds b) .6 seconds to 1.8 seconds c) .5 seconds to 2.1 seconds d) .4 seconds to 1.9 seconds

OpenStudy (callisto):

Can you try to solve it first?

OpenStudy (anonymous):

ok

Parth (parthkohli):

Oh wait this is easy

Parth (parthkohli):

It's just a quadratic inequality

Parth (parthkohli):

I dunno how to solve quadratic inequalities :\

OpenStudy (anonymous):

me either

OpenStudy (callisto):

Alright... it's like this: \[\frac{-(-6) -\sqrt{(-6)^2 - 4(4)(1)}}{2(4)} \le x \le \frac{-(-6) -\sqrt{(-6)^2 + 4(4)(1)}}{2(4)} \] Basically, it's just using the quadratic formula

OpenStudy (callisto):

Simplify it please.

OpenStudy (anonymous):

i got d .4 to 1.9 secs ???

OpenStudy (callisto):

 Nope, check it again!